A boat can travel 45 mph in still water. If it travels 200 miles with the current in the same length of time it travels 160 against the current,what is the speed of the current?
Answer by mananth(16946)

A boat can travel 45 mph in still water. If it travels 200 miles with the current in the same length of time it travels 160 against the current,what is the speed of the current?
...
let current speed be x mph
speed in still water = 45 mph
..
Upstream speed = 45-x mph
downstream speed = 45+x mph
..
time = Distance /speed
160/45-x = 200/45+x
160(45+x)=200(45-x)
7200+160x=9000-200x
200x+160x=9000-7200
360x=1800
x= 1800/360
x=5 mph the speed of current
...
m.ananth@hotmail.ca